On January 1, 2016, Howard, Inc. granted to a key executive a fixed compensatory share option plan for 1,000 shares of $4 par common stock for $30 a share. The fair value per option on that date was $14. The service period extended through December 31, 2017.

Refer to Exhibit 15-3. What entry, if any, was required on December 31, 2016?

no entry was necessary

Compensation Expense            7,000
Paid-in Capital Share Options                    7,000

Compensation Expense            6,000
Paid-in Capital Share Options                    6,000

Compensation Expense            9,000
Deferred Compensation                             9,000
Answer #1
Total Fair value of options = 1000*14 = 14000
The Compensation Expense is to be recognized equally over 2016 and 2017
The Compensation Expense for 2016 = 14000/2 = $7000
The entry required on December 31, 2016 is:
Compensation Expense            7,000
    Paid-in Capital Share Options                    7,000
Option B is correct